# Conceptualization, Data Curation

### v2

@@ 4.3 Auto-encoders
There are usually multiple operating modes in a chemical process, and fault detection models that are built based on one mode may have high false alarm rates. For multimode process monitoring and handling unseen modes a self-adaptive deep learning method based on local adaptive standardization and variational auto-encoder bidirectional long short-term memory (LAS-VB) is introduced in[97], in this research unstable tendency in the local moving window is detected. This method is not efficient for very slow-drif faults. Combining autoencoders and clustering algorithms for fault detection is a powerful tool, especially for dealing with unlabeled data. Zheng et al. [132], used a convolutional stacked autoencoder (SAE) network for feature extraction. Feature visualization is done by the t-distributed stochastic neighbour embedding (t-SNE) algorithm and DBSCAN clustering. The implemented method could detect Faults 1, 2, 4, 6, 7, 11, 13, 14, 17, 19 and 20 with average accuracy of 93%. Clustering is done unsupervised, however, a knowledge-based cluster annotation is used for labelling. Chadha et al. [11] also used a convolutional autoencoder for handling unlabeled data. The latent features are split into discriminative and reconstructive, and an auxiliary loss based on k-means clustering is used for discriminatory latent variables. The classification average accuracy for 3, 9, 15, and 21 faults which are very hard to detect is 50.04%. In [4], a source-aware autoencoder (SAAE) for detecting unseen faults and handling imbalanced data sets is introduced that uses Bidirectional long short-term memory (BiLSTM) and residual neural network (ResNet) to deal with randomness. The training can be done by using just normal data, and then fine-tuning by using emerged faulty samples. The proposed method detects unseen faults with an accuracy of 93.8%, moreover, the average accuracy of fault detection for different fault-to-normal ratios is 91.8%. In the balanced condition, the average accuracy is 94.3%. As usually there are not enough labelled samples, semi-supervised learning is developed that uses unlabeled data to help with the limited labelled data during model training. A consistency regularization autoencoder (CRAE) framework is introduced in [53], which captures correlations from labelled and unlabeled data. Previous samples are augmented with the current samples to increase the fault detection speed and accuracy. By just having the labels of just 20% of the dataset, the average fault detection accuracy of all the faults in the TE dataset except 3, 9, and 15 is 92.95%.
Peng et al. [68], introduced a non-revisiting genetic cost-sensitive sparse autoencoder(NrGCS-SAE), for solving the imbalanced fault diagnosis problem. The genetic algorithm optimizes class weights while the non-revisiting strategy increases exploration ability. This method could detect faults in the unbalanced TE dataset with an average accuracy of 92.02% while the normal to-fault ratio was 10:1. Convolutional autoencoders are powerful feature extraction tools, however, the pertaining phase is done unsupervised, to improve the extracted features and make sure they contain internal spatial information, supervised convolutional autoencoder (SCAE) is introduced [110]. The average accuracy of this method on the TE dataset is 96.19%, and it could reach an accuracy of about 95% in only two epochs. One-dimension residual convolutional auto-encoder (1DRCAE) is a model introduced in [111], for unsupervised feature extraction, where two residual learning blocks are embedded between the encoder and decoder to improve the speed and accuracy. The fault detection rate of the model on the TEP dataset on all faults except 3, 9, and 15 is 90.8%. To utilize unlabeled data more efficiently, Zhang et al. [126], combined LSTM for temporal feature extraction and ladder autoencoder (LAE) for semi-supervised learning. LSTM-LAE achieves the interpretability to extract fault-relevant process variables. By just having the labels of 11% of the dataset, this model achieved the average accuracy of 95.9% for all faults except 3,13, 15 and 16. Stacked spare-denoising autoencoder (SSDAE)-Softmax[46], is introduced for robust fault detection in complex industrial processes. The state transition algorithm (STA) is a global optimization method that is used for hyperparameter optimization. The fault detection accuracy of the model on the TEP reaches the average of 96.6% for faults 1, 2, 4, 5, 7, 13, and 17.
@@
Conventional autoencoder algorithms usually focus on the reconstruction error and neglect lots of useful information in the latent space for FDD. To address this problem one-dimension convolutional adversarial autoencoder (1DAAE) is introduced in [89], which uses the adversarial thought for clustering the latent variables to a prior distribution. Both reconstruction error and latent variable distribution are considered in the training of the model. The mean accuracy for all 21 faults of TEP is 86%. To monitor the nonlinear process and capture the process dynamics, a novel dynamic-inner convolutional autoencoder (DiCAE) is introduced in [125], that integrates a vector autoregressive model into a 1-dimensional convolutional autoencoder. The average FDR of 96.8% is achieved on the TEP for all the classes except 3, 9 and 15. Jang et al. [36], combined variational autoencoder and a generative adversarial network proposing an adversarial autoencoder (AAE), to generate features that follow the designed prior distribution. These features improve the reliability and stability of FDD algorithms. The average FDR for all 21 faults of TEP is 78.55%. This method decreases the fault detection delay and can detect faults faster compared to VAE. To reduce the irrelevant information with the faults during the feature extraction, a neuron-grouped stacked autoencoder (CG-SAE) is introduced in [62] hierarchical fault-related feature representation. In this method, hidden neurons are divided into different groups representing the data categories, to extract category-related features. This method achieved the average FDR of 96.02% for all faults except 3, 9 and 15. Li et al. [43], introduced a multimode feature fusion algorithm that selects the most relevant variables for each fault using the minimum redundancy-maximum relevance (mRMR) method, and then uses a stacked autoencoder for feature extraction. The average FDR for all faults except 3,9,15 is 88.04%, which is improved compared to SAE with the FDR of 76.77%. To reduce the noise in high-dimensional process signals, a one-dimensional convolutional auto-encoder (1D-CAE) is introduced in [14], This model does the feature extraction in a hierarchical structure, and this model outperforms SVM, DBN and SDAE. Jin et al. [37], used neurons with wavelet activation function (wavelons) and exponential linear units (ELU), to make a wavelons-constructed autoencoder-based deep neural network (WA-DNN), the average FDR for all 21 faults is 95.86%.
@@
Traditional fault detection techniques do not take quality indicators into account, they only look at process data. A mixing stacked auto-encoder, consisting of a nonlinear encoder and a linear decoder, is introduced in [102] to extract features of process variables and quality indicators while keeping feature extraction and model construction separate. Prior knowledge assists in enhancing the predictive accuracy of quality indicators and makes models easier to interpret in order to obtain more dependable monitoring performance. Stacked denoising auto-encoders (SDAE), are used to extract nonlinear latent variables and features from noisy data, however, the dynamic relationships are neglected. To consider them, a recursive stacked denoising auto-encoder (RSDAE) is proposed in [135]. For extracting more effective features from time domain process signals, a convolutional long short-term memory autoencoder (CLSTM-AE) is introduced in [112], where a residual block is used to select key information and improve the accuracy and reduce gradient disappearance and gradient explosion in the training phase. FDR for all the faults of TEP is 91%. Nonlocal and global preserving stacked autoencoder (NLGPSAE) [105], can extract structure-related features from nonlinear process data by using a regularized objective function that can control the reconstructed space. The average FDR for all the faults of TEP considering $T^{2}$ and SPE statistics are 94.8% and 91.8% respectively.
@@
To consider geometrical information from process signals during the feature extraction of SAEs in fault detection, manifold regularized stacked autoencoders (MRSAE) are introduced in [114] that considers the neighbourhood information. MRSAE does not need a large amount of labelled data. This model is superior in detecting certain types of faults like 5, 16, and 19, and the average FDR for 21 faults of TEP is 91.42%. To achieve interpretable analysis along the fault detection task, a novel orthogonal self-attentive variational autoencoder (OSAVA) model is introduced in [10]. Orthogonal attention extracts the correlations between variables and temporal dependency of different timesteps. The self-attention mechanisms utilized in the VSAE consider information from all timesteps to improve the performance. This model has a low detection delay an average of 24.2 minutes, the average FDR is 94.7%. To achieve effective and interpretable features for condition monitoring, a slow feature analysis-aided autoencoder (SFA-AE) is introduced in [40], which extracts deep slow variation patterns from the extracted features of AE. These slow patterns can show abnormal states. The approach uses a neural network consisting of a process variable selection (PVS) block, followed by an encoder-decoder structure and a deep slow feature analysis (DSFA) section. The PVS block utilizes 1D convolutional layers to generate an attention map that represents feature importance for selecting crucial process variables. Next, the encoder and decoder senctions use convolutional LSTM and deconvolutional LSTM layers respectively to learn spatiotemporal relationships of the input variables. The average FDR for 21 faults of TEP is 94.87%. To handle nonlinearity, non-Gaussian features, and dynamics in process monitoring of complex systems, independent component analysis (ICA) and adversarial autoencoder (AAE) are fused in [16].
4.4 Generative Adversarial Network (GAN)
@@
Generative Adversarial Networks (GAN) are deep learning models utilized for dealing with imbalanced data as they are generative algorithms that can increase the number of samples and improve data-driven algorithm accuracy by providing more data. GANs consist of two networks, a generator model that tries to generate synthetic samples and a discriminator model that discovers if a sample is generated by the generator or not[24]. A schematic of a GAN is shown in 8.
Figure 8: A schematic representation a GAN[3]
@@
For solving the imbalance fault diagnosis problem an enhanced auxiliary classifier generative adversarial network (EACGAN) is introduced [67]. In this model, the boundary-seeking loss is used for stabilizing the training process, with the normal to-fault ratio of 10:1 and considering all faults in the TEP dataset except 3, 9 and 15, the accuracy of the model is 77.01%. High-efficiency GAN model (HGAN) [74] is a model that integrates Wasserstein GAN and Auxiliary Classifier GAN to improve the stability and efficiency of the training process. This model is not just limited to generating data but directly can be used for fault classification. For the balanced TEP dataset, considering all faults the accuracy is 86.6%, while by the normal to fault samples ratio of 5:1, the accuracy is 82.0%. Sparse semi-supervised GAN (SSGAN) [47] is a model that uses unlabeled and labelled data for training the model. In this model, the discriminator is changed into a multi-dimensional classifier and the loss function for each labelled, unlabeled and generated type of data is defined separately, and the total loss is used in optimization. Lomov et al. [48], studied different recurrent and convolutional networks and introduced a novel temporal CNN1D2D architecture. A GAN is used to enrich the training data samples. This method performs well in the detection of 3, 9, and 15 faults, which are very challenging.
@@ 4.5 Deep Belief Network (DBN)
Deep Belief Networks are primarily used for unsupervised learning tasks, especially for feature learning, they are made of multilayer Restricted Boltzmann machines (RBMs), which are energy-based probabilistic models. DBN consist of visible and hidden layers, the visible layer represents the input data, while the hidden layers capture abstract and complex features. Each layer is fully connected to the adjacent layers, but there are no connections within a single layer. A schematic of DBN is shown in 9 .
Figure 9: A schematic representation of DBN [93]
@@
One of the first researches that used DBN for chemical plants FDD is [99], unlabeled data is used to pre-trian the DBN, and then labelled data is used for training the DNN using backpropagation. The average classification accuracy is 80.5%. Using deep learning techniques for layer-by-layer feature extraction may lead to the weakening or vanishing of the fluctuations that reflect the fault information. To address this issue, in [113], the support vector data description (SVDD) and moving average filter technologies are utilized, moreover, in the structure of DBN, the traditional Restricted Boltzmann Machine (RBM) is replaced with the Gaussian Restricted Boltzmann Machine (GRBM). To reduce overfitting and improve the generalization of DBN, dropout is used in DBN [93]. Dropout randomly disables a certain portion of units in the training process. By using this method the average F1-score for all faults of TEP is 0.81. By using deep learning for FDD, important information in the raw data might be filtered during feature compression, to solve this problem an extended deep belief network (EDBN) is proposed in [92], where each extended restricted Boltzmann machine (ERBM) receives raw data and hidden features as inputs during the pre-training phase. The dynamic properties of process data are then taken into account while building a dynamic EDBN-based fault classifier. The average FDR of all TEP faults except 3,9,15 using EDBN is 94.31%, which is slightly better than DBN 93.89%. The existing noise level in industrial environments affects the performance of deep learning FDD models. To reduce the noise effects, a penalty factor is added to the network structure, avoiding the local optimal situation of a DBN and increasing FDD accuracy, besides an adaptive lifting wavelet (ALW) that reduces the noise level leading to ALW-DBN model[106]. The average FDR of TEP for all faults is 97.33%. However, The ALW-DBN model’s training takes longer because of the iterative penalty optimisation algorithm. Additionally, the DBN method has obvious drawbacks when attempting to extract the time information from chemical data in comparison to the DCNN method. Lui et al. [44], introduced an index for measuring fault information of high dimensional features and finding deep highly-sensitive features (DHSFs) from the extracted features of a DBN. In the mentioned research Gauss– Bernoulli Restricted Boltzmann Machine (GRBM) is used to replace the traditional binary Restricted Boltzmann Machine (RBM), to reduce information loss. Finally, Euclidean distance (ED) is used to calculate the distance between normal and fault samples and the moving average (MA) to eliminate the effect of noise. The average FDR of the proposed DHSF-DBN method is 88.67%. Yang et al. [104] proposed a pairwise graph regularized deep belief network (PG-DBN) model for FDD. Two different graph constraints are considered on the hidden layer of the RBM, preserving the feature manifold structure in the same class of the data and penalizing the feature manifold structure in different classes of the data, respectively. The average classification accuracy of PG-DBN, G-DBN, CNN and DBN which are studied in this paper is 93.87%, 91.98%, 89.18% and 83.1% for all 21 faults of TEP.

### v3

@@ 3.5 Auto-encoders
There are usually multiple operating modes in a chemical process, and fault detection models that are built based on one mode may have high false alarm rates. For multimode process monitoring and handling unseen modes, a self-adaptive deep learning method based on local adaptive standardization and variational auto-encoder bidirectional long short-term memory (LAS-VB) is introduced in Wu and Zhao (2020). In this research, unstable tendencies in the local moving window are detected. This method is not efficient for very slow-drift faults. Combining autoencoders and clustering algorithms for fault detection is a powerful tool, especially for dealing with unlabeled data. Zheng et al. Zheng and Zhao (2020) used a convolutional stacked autoencoder (SAE) network for feature extraction. Feature visualization is done by the t-distributed stochastic neighbour embedding (t-SNE) algorithm and DBSCAN clustering. The implemented method could detect Faults 1, 2, 4, 6, 7, 11, 13, 14, 17, 19, and 20 with an average accuracy of 93%. Clustering is done unsupervised, and a knowledge-based cluster annotation is used for labeling.
@@
Convolutional autoencoders are powerful feature extraction tools; however, the pertaining phase is done unsupervised. To improve the extracted features and make sure they contain internal spatial information, a supervised convolutional autoencoder (SCAE) is introduced Yu et al. (2022a). The average accuracy of this method on the TE dataset is 96.19%, and it could reach an accuracy of about 95% in only two epochs. One-dimension residual convolutional auto-encoder (1DRCAE) is a model introduced in Yu and Liu (2022), for unsupervised feature extraction, where two residual learning blocks are embedded between the encoder and decoder to improve the speed and accuracy. The fault detection rate of the model on the TEP dataset on all faults except 3, 9, and 15 is 90.8%. Stacked spare-denoising autoencoder (SSDAE)-SoftmaxLiu et al. (2021b), is introduced for robust fault detection in complex industrial processes. The state transition algorithm (STA) is a global optimization method that is used for hyperparameter optimization. The fault detection accuracy of the model on the TEP reaches an average of 96.6% for faults 1, 2, 4, 5, 7, 13, and 17. Conventional autoencoder algorithms typically prioritize the reconstruction error, overlooking valuable information in the latent space for FDD. To address this problem, a one-dimensional convolutional adversarial autoencoder (1DAAE) is introduced in Wang et al. (2023a), which uses the adversarial thought for clustering the latent variables to a prior distribution. Both reconstruction error and latent variable distribution are considered in the training of the model. The mean accuracy for all 21 faults of TEP is 86%. To monitor the nonlinear process and capture the process dynamics, a novel dynamic-inner convolutional autoencoder (DiCAE) is introduced in Zhang and Qiu (2022a), integrating a vector autoregressive model into a 1-dimensional convolutional autoencoder. The average FDR of 96.8% is achieved on the TEP for all the classes except 3, 9, and 15. Jang et al. Jang et al. (2021), combined variational autoencoder and a generative adversarial network proposing an adversarial autoencoder (AAE), to generate features that follow the designed prior distribution. These features improve the reliability and stability of FDD algorithms. The average FDR for all 21 faults of TEP is 78.55%. This method decreases the fault detection delay and can detect faults faster compared to VAE. To reduce the irrelevant information with the faults during feature extraction, a neuron-grouped stacked autoencoder (CG-SAE) is introduced in Pan et al. (2021) hierarchical fault-related feature representation. In this method, hidden neurons are divided into different groups representing the data categories to extract category-related features. This method achieved the average FDR of 96.02% for all faults except 3, 9, and 15. Li et al. Li et al. (2020b) introduced a multimode feature fusion algorithm that selects the most relevant variables for each fault using the minimum redundancy-maximum relevance (mRMR) method, and then uses a stacked autoencoder for feature extraction. The average FDR for all faults except 3,9, and 15 is 88.04%, which is improved compared to SAE with the FDR of 76.77%. To reduce the noise effect in high-dimensional process signals, a one-dimensional convolutional auto-encoder (1D-CAE) is introduced in Chen et al. (2020).. This model performs feature extraction in a hierarchical structure, outperforming SVM, DBN, and SDAE. Jin et al. Jin et al. (2018) utilized neurons with a wavelet activation function (Wavelons) and exponential linear units (ELU) to construct a Wavelons-constructed autoencoder-based deep neural network (WA-DNN), achieving an average FDR of 95.86% across all 21 faults.
@@
Traditional fault detection techniques do not take quality indicators into account; they only look at process data. A mixing stacked auto-encoder, consisting of a nonlinear encoder and a linear decoder, is introduced in Yan and Yan (2019) to extract features of process variables and quality indicators while keeping feature extraction and model construction separate. Prior knowledge enhances the predictive accuracy of quality indicators and makes models easier to interpret, leading to more dependable monitoring performance. Stacked denoising auto-encoders (SDAE), are used to extract nonlinear latent variables and features from noisy data, however, the dynamic relationships are neglected. To consider them, a recursive stacked denoising auto-encoder (RSDAE) is proposed in Zhu et al. (2020). For extracting more effective features from time domain process signals, a convolutional long short-term memory autoencoder (CLSTM-AE) is introduced in Yu et al. (2020), where a residual block is used to select key information and improve the accuracy, and reduce gradient disappearance and gradient explosion in the training phase. FDR for all the faults of TEP is 91%. Nonlocal and global preserving stacked autoencoder (NLGPSAE) Yang and Wang (2023) can extract structure-related features from nonlinear process data by using a regularized objective function that can control the reconstructed space. The average FDR for all the faults of TEP, considering $T^{2}$ and SPE statistics, is 94.8% and 91.8%, respectively.
@@
To consider geometrical information from process signals during the feature extraction of SAEs in fault detection, manifold regularized stacked autoencoders (MRSAE) are introduced in Yu and Zhang (2020) that consider the neighbourhood information. MRSAE does not need a large amount of labeled data. This model is superior in detecting certain types of faults like 5, 16, and 19, and the average FDR for 21 faults of TEP is 91.42%. To achieve interpretable analysis along the fault detection task, a novel orthogonal self-attentive variational autoencoder (OSAVA) model is introduced in Bi and Zhao (2021). Orthogonal attention extracts the correlations between variables and the temporal dependency of different time steps. The self-attention mechanisms utilized in the VSAE consider information from all time steps to improve the performance. This model has a low detection delay, an average of 24.2 minutes, and the average FDR is 94.7%. To achieve effective and interpretable features for condition monitoring, a slow feature analysis-aided autoencoder (SFA-AE) is introduced in Li et al. (2021), which extracts deep slow variation patterns from the extracted features of AE. These slow patterns can show abnormal states. The approach uses a neural network consisting of a process variable selection (PVS) block, followed by an encoder-decoder structure and a deep slow feature analysis (DSFA) section. The PVS block utilizes 1D convolutional layers to generate an attention map that represents feature importance for selecting crucial process variables. Next, the encoder and decoder sections use convolutional LSTM and deconvolutional LSTM layers, respectively, to learn spatiotemporal relationships of the input variables. The average FDR for 21 faults of TEP is 94.87%. To handle nonlinearity, non-Gaussian features, and dynamics in process monitoring of complex systems, independent component analysis (ICA) and adversarial autoencoder (AAE) are fused in Chen et al. (2025). To improve detection reliability, fault thresholds are determined using tail distribution features, achieving the average FDR of 99.52% for all the faults except 3, 9, and 15. Residual squeeze-and-excitation convolutional auto-encoder (RSECAE) is introduced in Yu et al. (2024) for effective feature extraction. This method proposes a hybrid structure combining convolution calculation, an auto-encoder, and multiple residual squeeze-and-excitation networks (RSENets), and utilizing maximum mean discrepancy (MMD) to reduce the distribution difference between the process signals and the learned features. This method is evaluated using TEP and a three-phase process (real industrial process), showing the ability of the unsupervised learning method in the absence of fault data. A differential recurrent autoencoder (DRAE) is proposed in Ji et al. (2025) to address the challenge of early fault detection in nonstationary industrial processes, where traditional methods struggle to capture transient, time-varying behaviors caused by stochastic disturbances and operational variability. By embedding differential operations into the latent layer of LSTM networks, DRAE enhances the forget gate’s responsiveness to short-term nonstationary features while preserving long-term dependencies. The model leverages an unsupervised encoder-decoder structure, using reconstruction error as a fault indicator, effectively distinguishing early-stage faults from inherent nonstationary trends. Validation on synthetic data, a real-world chemical plant (industrial continuous catalytic reforming heat exchange unit), and TEP achieving the average FDR of 92.34% for faults 1 to 14 except 3 and 9, demonstrates that DRAE significantly improves fault detection rates and reduces detection delays compared to conventional methods.
Autoencoder-based methods are highly effective for unsupervised feature learning and dimensionality reduction, making them suitable when labeled fault data are scarce. Temporal integration (e.g., CLSTM-AE, OSAVA) boosts performance by capturing time-dependent variations in latent space, while supervised variations (e.g., SCAE) accelerate convergence and improve discriminative ability. Incorporating sparsity, manifold regularization, or adversarial learning helps enhance robustness against noise and improve separation of fault classes in latent space. However, autoencoders generally underperform in detecting extremely subtle or slowly developing faults unless paired with temporal models or variable selection mechanisms. Their success also heavily depends on the quality of normal training data to avoid false alarms. A summary of these methods is shown in 5.
Table 5: Summary of Autoencoder Models. See Table 2 for fault types.
Model Reference Specification Faults Information Average Accuracy
WA-DNN Jin et al. (2018) Using wavelet activation function in the autoencoder All 95.86%
CLSTM-AE Yu et al. (2020) Convolutional LSTM memory autoencoder (CLSTM-AE) and residual block for increasing the accuracy and making the training process more robust All 91%
MRSAE Yu and Zhang (2020) Manifold regularized stacked autoencoders (MRSAE) considering neighbourhood information needs a few labeled data All 91.42%
DiCAE Zhang and Qiu (2022a) integrating vector autoregressive model and 1-dimensional convolutional autoencoder for capturing the process dynamics All faults except 3, 9, and 15 96.80%
CG-SAE Pan et al. (2021) Neuron-grouped stacked autoencoder (CG-SAE) as a hierarchical feature extractor for different groups of faults All faults except 3, 9, and 15 96.02%
OSAVA Bi and Zhao (2021) Orthogonal self-attentive variational autoencoder (OSAVA) to increase the interpretability with a low detection delay All faults except 15, 16 94.70%
SFA-AE Li et al. (2021) Slow feature analysis-aided autoencoder (SFA-AE) to deal with slow variations in the data All faults except 15, 17 94.87%
LSTM-SAE Agarwal et al. (2022) Combining LSTM and Supervised Autoencoder, improving sensitivity by different models for different faults All 93.23%
VAE + RNN Cheng et al. (2019) Variational recurrent autoencoder combines RNN and VAE to handle nonlinearity All faults except 3,9,15 96.30%
SCAE Yu et al. (2022a) Supervised convolutional autoencoder (SCAE) for better feature extraction, high accuracy in a few epochs All faults except 3,9,15,16 96.19%
1DAAE Wang et al. (2023a) Clustering the latent variables, considering both reconstruction error and latent space information All 86%
NLGPSAE Yang and Wang (2023) Nonlocal and global preserving stacked autoencoder (NLGPSAE) using a regularized objective function and extracting structure-related features All 94.80%
ICA–AAE Chen et al. (2025) Handle nonlinearity, non-Gaussian features All faults except 3,9,15 99.52%
DRAE Ji et al. (2025) Captures transient, time-varying behaviors. Early-stage fault detection. Also a real-world chemical plant faults 1 to 14 except 3 and 9 92.34%
3.6 Generative Models
@@
Generative models are machine learning algorithms that learn patterns and distributions of the data to generate new data, including models such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and diffusion models. GANs are deep learning models utilized for dealing with imbalanced data, as they are generative algorithms that can increase the number of samples and improve data-driven algorithm accuracy by providing more data. GANs consist of two networks, a generator model that tries to generate synthetic samples and a discriminator model that discovers if a sample is generated by the generator or not Goodfellow et al. (2020).
@@
For solving the imbalance fault diagnosis problem, an enhanced auxiliary classifier generative adversarial network (EACGAN) is introduced Peng et al. (2020a). In this model, the boundary-seeking loss is used for stabilizing the training process, with the normal to-fault ratio of 10:1 and considering all faults in the TEP dataset except 3, 9, and 15, the accuracy of the model is 77.01%. High-efficiency GAN model (HGAN) Qin and Zhao (2022b) is a model that integrates Wasserstein GAN and Auxiliary Classifier GAN to improve the stability and efficiency of the training process. This model is not just limited to generating data, but can also be used directly for fault classification. For the balanced TEP dataset, considering all faults, the accuracy is 86.6%, while by the normal to fault samples ratio of 5:1, the accuracy is 82.0%. Sparse semi-supervised GAN (SSGAN) LIU et al. (2020) is a model that uses unlabeled and labeled data for training the model. In this model, the discriminator is transformed into a multi-dimensional classifier. The loss functions for labeled, unlabeled, and generated data types are defined separately, and the total loss is used in optimization. Lomov et al. Lomov et al. (2021) studied different recurrent and convolutional networks and introduced a novel temporal CNN1D2D architecture. A GAN is used to enrich the training data samples. This method performs well in the detection of 3, 9, and 15 faults, which are very challenging. Shirshahi et al. Shirshahi et al. (2025) introduced a novel architecture for detecting similar patterns and preventing alarm floods in industrial systems. First, process variables and alarm sequences are jointly represented and pre-processed; these fused data are then mapped into a latent space by a VAE. A GAN measures similarity between latent patterns, and multi-sensor information fusion aggregates similarity indices to identify the root cause of the faults. By incorporating probability density functions of the extracted features, the method mitigates the influence of random noise. The online implementation supports early fault detection and rapid operator decision-making, demonstrating efficacy on the TEP simulator. Other generative models, such as diffusion models, can be used for data generation to enhance FDD in industrial systems, eliminating the need for adversarial training. In Li et al. (2025), an improved supervised contrastive learning (ISCL) with a denoising diffusion probabilistic model (DDPM) is used to address the challenge of overlapped normal and abnormal data distributions.
In summary, generative approaches can address imbalanced and scarce fault data challenges by synthesizing realistic samples for minority fault classes. Augmenting training sets via GANs can significantly improve classification balance, especially for rare faults. VAEs and hybrid VAE–GAN models provide structured latent spaces, assisting in both anomaly detection and interpretability. Diffusion models, though less explored, show promise for stable, high-quality data generation without adversarial training challenges. Yet, generative models often require careful tuning to prevent mode collapse or generation of unrealistic samples, and their benefit is less noticeable when the original dataset is already large and balanced.
@@ 3.7 Deep Belief Network (DBN)
Deep Belief Networks are primarily used for unsupervised learning tasks, especially for feature learning. They are made of multilayer Restricted Boltzmann machines (RBMs), which are energy-based probabilistic models. DBN consists of visible and hidden layers; the visible layer represents the input data, while the hidden layers capture abstract and complex features. Each layer is fully connected to the adjacent layers, but there are no connections within a single layer.
@@
One of the first studies that used DBN for chemical plants FDD is Xie and Bai (2015), unlabeled data is used to pre-train the DBN, and then labeled data is used for training the DNN using backpropagation. The average classification accuracy is 80.5%. Using deep learning techniques for layer-by-layer feature extraction may lead to the weakening or vanishing of the fluctuations that reflect the fault information. To address this issue, in Yu and Yan (2018), the support vector data description (SVDD) and moving average filter technologies are utilized; moreover, in the structure of DBN, the traditional Restricted Boltzmann Machine (RBM) is replaced with the Gaussian Restricted Boltzmann Machine (GRBM). To reduce overfitting and improve the generalization of DBN, dropout is used in DBN Wei and Weng (2020). Dropout randomly disables a certain portion of units in the training process. By using this method, the average F1-score for all faults of TEP is 0.81. By using deep learning for FDD, important information in the raw data might be filtered during feature compression. To solve this problem, an extended deep belief network (EDBN) is proposed in Wang et al. (2020), where each extended restricted Boltzmann machine (ERBM) receives raw data and hidden features as inputs during the pre-training phase. The dynamic properties of process data are then taken into account while building a dynamic EDBN-based fault classifier. The average FDR of all TEP faults except 3, 9, 15 using EDBN is 94.31%, which is slightly better than DBN 93.89%. The existing noise level in industrial environments affects the performance of deep learning FDD models. To reduce the noise effects, a penalty factor is added to the network structure, avoiding the local optimal situation of a DBN and increasing FDD accuracy, besides an adaptive lifting wavelet (ALW) that reduces the noise level, leading to the ALW-DBN modelYao et al. (2021). The average FDR of TEP for all faults is 97.33%. However, the ALW-DBN model’s training takes longer because of the iterative penalty optimization algorithm. Additionally, the DBN method has obvious drawbacks when attempting to extract the time information from chemical data in comparison to the DCNN method. Lui et al. Liu et al. (2021a) introduced an index for measuring fault information of high-dimensional features and finding deep highly-sensitive features (DHSFs) from the extracted features of a DBN. In the mentioned research, the Gauss-Bernoulli Restricted Boltzmann Machine (GRBM) is used to replace the traditional binary Restricted Boltzmann Machine (RBM) to reduce information loss. Finally, Euclidean distance (ED) is used to calculate the distance between normal and fault samples, and the moving average (MA) to eliminate the effect of noise. The average FDR of the proposed DHSF-DBN method is 88.67%. Yang et al. Yang et al. (2021) proposed a pairwise graph regularized deep belief network (PG-DBN) model for FDD. Two different graph constraints are considered on the hidden layer of the RBM, preserving the feature manifold structure in the same class of the data and penalizing the feature manifold structure in different classes of the data, respectively. The average classification accuracy of PG-DBN, G-DBN, CNN, and DBN, which are studied in this paper, is 93.87%, 91.98%, 89.18%, and 83.1% for all 21 faults of TEP.
DBNs demonstrate competitive performance when combined with enhancements such as graph regularization (PG-DBN) or noise reduction (ALW-DBN). These methods handle nonlinearities well and can be trained with limited labeled data, making them attractive for industrial contexts where annotation is expensive. However, DBNs are computationally heavier than CNN or AE alternatives and often lag in capturing long temporal dependencies, limiting their performance in processes with slow or gradual fault progression. Recent DBN variants integrate temporal features explicitly, but this area remains less developed compared to RNN or CNN hybrids.